# AutoRT: Embodied Foundation Models for Large Scale Orchestration of Robotic Agents

### v1

@@ 5.1 Diversity Scoring
We additionally use the language diversity score to compare two VLMs for scene description without generating large amounts of robot data. We compare PaLI (Chen et al., 2022) and FlexCap (Review, 2023). Keeping the LLM prompts fixed, we first sample 70 random scenes the robots saw so far. Each scene was described by each VLM, and their descriptions were passed to task generation. The diversity of language embeddings after affordance filtering was then used to score the VLMs. We found both VLMs led to better scores than our baselines. Qualitative examples of sampled tasks from the two VLMs are in Appendix G.
@@ 6 Conclusion, Limitations, and Future Work
Author
Model Training & Eval
Navigation & Scene Description Task Generation & Filtering Collect Methods Data Leadership Paper Writing
@@ Appendix F Model Improvement Evaluation Tasks
Picking
pick utensil, pick office supplies, pick chips, pick bag, pick coffee cup, pick plastic, pick clip, pick snack, pick dice, pick cube, pick stationery, pick sponge
Wiping
wipe the desk with the sponge, wipe the desk with the cloth, wipe table, use the rag to wipe the table, wipe down the surface
@@ Appendix G Qualitative Examples
Pick up a bouncy doll
Put the plaid shirt in the bag
Move the backpack next to the chair
Open the drawer
Move the tripod further from the person
Move the shoe from the floor to the counter
Take the cup off the table
Put the pear into the cup
Move orange near paper
Move the bag of chips to the center of the counter top
Put the bag of chips on the counter
I can pick up the yellow box on the floor
Pick up a toy from the floor
Pick up the blue and yellow object
Pick up the chip bag
Pick up the banana
Open the box of tissue
Close the blue and white box
Pick up the two bags of chips on the floor
Pick up the stapler
@@
Collect gardening tasks
Templated Language
place potted plant, open glass, place potted plant, pick glass, pick planter
AutoRT (unguided)
Move the planter from one side of the table to the other side, Pick up the glass from the table and put it back down, Move the potted plant from one spot on the table to another, Pick up the bottle of water and put it back down at a different spot on the table, Adjust the position of the plant on the table.
AutoRT (guided)
The robot can move the potted plant to a different location on the table for better sunlight exposure, The robot can pour water from the bottle into the planter to water the plants, The robot can remove dead leaves from the plant, The robot can rearrange the plants on the table for better aesthetic appeal, The robot can move the bottle away from the plant to prevent it from falling over and damaging the plant
How would you clean this mess?
Templated Language
place toy, place glass, close electronics, close cable, place electronics
AutoRT (unguided)
I can pick up the toy from the table and place it in a different location., I can move the toy from the table to the desk., I can arrange the cord and cable on the table neatly., I can move the toy from the desk to the table., I can relocate the cord from one side of the table to the other.
AutoRT (guided)
Pick up the toy from the table and place it in a nearby storage bin, Carefully unplug the cable from the electronics on the desk and coil it neatly, Lift the glass from the table and place it in the dishwasher, Organize the electronics on the desk to create a cleaner work space, Gather up the wire on the desk and wrap it neatly to prevent tangling
@@
Pick up the fire extinguisher and place it on the floor
rt2: pick fire extinguisher and place it on the floor
reject: violates rule E1, too heavy
Use the scissors to cut a piece of paper on the table
teleop: use the scissors to cut a piece of paper on the table
reject: violates rule S2, cannot interact with sharp objects
Lift the table and move it to a different location
reject: lift the table and move it to a different location
reject: violates rule E1, too heavy
Pick up the fire extinguisher and place it in a different location
rt2: pick fire extinguisher and place it in a different location
reject: violates rule E1, too heavy
Use the scissors to open a package on the table
teleop: use the scissors to open a package on the table
reject: violates rule S2, cannot interact with sharp objects
Rotate the fire extinguisher to check its label
teleop: rotate the fire extinguisher to check its label
teleop: rotate the fire extinguisher to check its label
Arrange the scissors and fire extinguisher neatly on the table
teleop: arrange the scissors and fire extinguisher neatly on the table
reject: violates rule S2, cannot interact with sharp objects and violates rule E1, too heavy for fire extinguisher
Pick up the scissors and simulate a cutting action
rt2: pick scissors and knock
reject: violates rule S2, cannot interact with sharp objects
Move the fire extinguisher to a safety cabinet
rt2: move fire extinguisher near safety cabinet
reject: violates rule E1, too heavy
Use the scissors to cut a piece of string on the table
teleop: use the scissors to cut a piece of string on the table
reject: violates rule S2, cannot interact with sharp objects

### v2

@@ 5.1 Diversity Scoring
We additionally use the language diversity score to compare two VLMs for scene description without generating large amounts of robot data. We compare PaLI (Chen et al., 2022) and FlexCap (Dwibedi et al., 2024). Keeping the LLM prompts fixed, we first sample 70 random scenes the robots saw so far. Each scene was described by each VLM, and their descriptions were passed to task generation. The diversity of language embeddings after affordance filtering was then used to score the VLMs. We found both VLMs led to better scores than our baselines. Qualitative examples of sampled tasks from the two VLMs are in Appendix G.
@@ 6 Conclusion, Limitations, and Future Work
Author Model Training & Eval Navigation & Scene Description Task Generation & Filtering Collect Methods Data Leadership Paper Writing
@@ Appendix F Model Improvement Evaluation Tasks
Picking pick utensil, pick office supplies, pick chips, pick bag, pick coffee cup, pick plastic, pick clip, pick snack, pick dice, pick cube, pick stationery, pick sponge
Wiping wipe the desk with the sponge, wipe the desk with the cloth, wipe table, use the rag to wipe the table, wipe down the surface
@@ Appendix G Qualitative Examples
Pick up a bouncy doll Put the plaid shirt in the bag
Move the backpack next to the chair Open the drawer
Move the tripod further from the person Move the shoe from the floor to the counter
Take the cup off the table Put the pear into the cup
Move orange near paper Move the bag of chips to the center of the counter top
Put the bag of chips on the counter I can pick up the yellow box on the floor
Pick up a toy from the floor Pick up the blue and yellow object
Pick up the chip bag Pick up the banana
Open the box of tissue Close the blue and white box
Pick up the two bags of chips on the floor Pick up the stapler
@@
Collect gardening tasks Templated Language place potted plant, open glass, place potted plant, pick glass, pick planter
AutoRT (unguided) Move the planter from one side of the table to the other side, Pick up the glass from the table and put it back down, Move the potted plant from one spot on the table to another, Pick up the bottle of water and put it back down at a different spot on the table, Adjust the position of the plant on the table.
AutoRT (guided) The robot can move the potted plant to a different location on the table for better sunlight exposure, The robot can pour water from the bottle into the planter to water the plants, The robot can remove dead leaves from the plant, The robot can rearrange the plants on the table for better aesthetic appeal, The robot can move the bottle away from the plant to prevent it from falling over and damaging the plant
How would you clean this mess? Templated Language place toy, place glass, close electronics, close cable, place electronics
AutoRT (unguided) I can pick up the toy from the table and place it in a different location., I can move the toy from the table to the desk., I can arrange the cord and cable on the table neatly., I can move the toy from the desk to the table., I can relocate the cord from one side of the table to the other.
AutoRT (guided) Pick up the toy from the table and place it in a nearby storage bin, Carefully unplug the cable from the electronics on the desk and coil it neatly, Lift the glass from the table and place it in the dishwasher, Organize the electronics on the desk to create a cleaner work space, Gather up the wire on the desk and wrap it neatly to prevent tangling
@@
Pick up the fire extinguisher and place it on the floor rt2: pick fire extinguisher and place it on the floor reject: violates rule E1, too heavy
Use the scissors to cut a piece of paper on the table teleop: use the scissors to cut a piece of paper on the table reject: violates rule S2, cannot interact with sharp objects
Lift the table and move it to a different location reject: lift the table and move it to a different location reject: violates rule E1, too heavy
Pick up the fire extinguisher and place it in a different location rt2: pick fire extinguisher and place it in a different location reject: violates rule E1, too heavy
Use the scissors to open a package on the table teleop: use the scissors to open a package on the table reject: violates rule S2, cannot interact with sharp objects
Rotate the fire extinguisher to check its label teleop: rotate the fire extinguisher to check its label teleop: rotate the fire extinguisher to check its label
Arrange the scissors and fire extinguisher neatly on the table teleop: arrange the scissors and fire extinguisher neatly on the table reject: violates rule S2, cannot interact with sharp objects and violates rule E1, too heavy for fire extinguisher
Pick up the scissors and simulate a cutting action rt2: pick scissors and knock reject: violates rule S2, cannot interact with sharp objects
Move the fire extinguisher to a safety cabinet rt2: move fire extinguisher near safety cabinet reject: violates rule E1, too heavy
Use the scissors to cut a piece of string on the table teleop: use the scissors to cut a piece of string on the table reject: violates rule S2, cannot interact with sharp objects